The mole fraction of an aqueous solution of manganese(II) iodate is 0.266. Calculate the molarity (in mol/L) of the manganese(II) iodate solution, if the density of the solution is 1.48 g mL-1
1
Expert's answer
2019-11-07T08:04:06-0500

Mn(IO3)2 = 405 g/mol.x 0.266 = 107.73 g

H2O= 18 g/mol x 0.734 = 13.212 g.

107.73+13.212 = 120.942 g.

V= 120.942/1.48 = 81.72 ml / 1000 = 0.08172 L.

Cm= n/V = 0.266 / 0.08172 = 3.255 M.